A web developer is creating a website. The page will have a frame along its top and left sides showing the name of the site and links to different parts of the site. These frames will take up one third of the computer screen. If the computer screen is 7 in by 11 in, what will the width of the frames be?
Answer by jorel555(1290)

The area of the screen is 7 x 11=77
Let n be the width of the frames. Then:
(7-n)(11-n)=(2x77)/3
231-54n+3nē=154
3nē-54n+77=0
Using the quadratic formula; we get n=16.439 and 1.561. Throwing out the higher number, we get the width of the frame to be 1.561 inches. ☺☺☺☺